Operation Summary for May 31, 2010 - June 2, 2010 |
Monday May 31 The day shift began with Operations monitoring the Tevatron (TeV) store 7860 (100.37E30) and the Recycler stash (162.3E10), with the Antiproton Source (Pbar) stacking (13.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the FTBF T-1006 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations disabled the T-1006 beam at 9:28 AM. Operations terminated store 7860 at 3:06 PM. Operations prepared D0 for an access that continued through to the end of the shift. Operations reported some Booster downtime due to an ion pump (IPRF14) tripping off. During the evening shift, D0 came out of access and Operations put the TeV into shot setup at 6:10 PM. Operations established store 7863 at 7:14 PM with an average initial luminosity (AIL) of 366.55E30. Operations reported more Booster downtime caused by IPRF14 tripping off. Tuesday June 1 The midnight shift began with Operations monitoring store 7863 (159.27E30) and the stash (138.5E10), with Pbar stacking (19.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported some downtime due to a Booster RF station (BRF15) LCW flow fault that caused the West Anode Power Supply (WAPS) to trip off; everything reset. During the day shift, Operations terminated store 7863 at 10:29 AM and then put the TeV into shot setup. Operations established store 7864 at 11:55 AM with an AIL of 355.3E30. Instrumentation technicians replaced a failed MI flying wire VME power supply. Operators prepared Switchyard enclosures C, D, and E for access. Pbar and MI-8 personnel conducted collimator studies. Booster IPRF14 tripped several times and caused some downtime. During the evening shift, Operations reported some downtime due to several Linac RF station (LRF5) trips. Wednesday June 2 The midnight shift began with Operations monitoring store 7864 (84.38E30) and the stash (271.2E10), with Pbar stacking (26.3E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, store 7864 quenched at 12:22 AM due to an antiproton abort kicker pre-fire. Several TeV cryo systems went down, with TeV sector A1 being hit the hardest. Recovery took the remainder of the shift. The abort kickers were put in condition mode until around 6 AM. Operations reported a swirl of nasty thunderstorms that swept through the area, causing some chaos: a MI power supply interlock loop trip, various RF system trips, and a MINOS Electrical Safety System trip. Operators prepared the TeV A-E enclosures and the Transfer Hall for access to do vacuum work. The Plans for Wednesday The plans for today are to recover the TeV cryo systems, conduct TeV studies, and then resume stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 06/02/2010 email Fermilab |
| Security, Privacy, Legal | |